# An Optimal Design Framework for Lasso Sign Recovery

### v2

@@ 2 Exact Local Optimality Criteria
Remark 2 The probability of support recovery, $P(\hat{\mathcal{A}}=\mathcal{A}\,|\,\textbf{F},\,\boldsymbol{\beta})$, requires consideration of all possible $2^{k}$ sign vectors that have the same 00 elements as $\boldsymbol{z}$ but alternative ±$\pm 1$ elements indexed by $\mathcal{A}$. Defining $\mathcal{Z}_{\mathcal{A}}$ as the set of all such sign vectors, we have $P(\hat{\mathcal{A}}=\mathcal{A}\,|\,\textbf{F},\,\boldsymbol{\beta})=\sum_{\tilde{\boldsymbol{z}}\in\mathcal{Z}_{\mathcal{A}}}P(\hat{\boldsymbol{z}}=\tilde{\boldsymbol{z}}\,|\,\textbf{F},\,\boldsymbol{\beta})$. Each $P(\hat{\boldsymbol{z}}=\tilde{\boldsymbol{z}}\,|\,\textbf{F},\,\boldsymbol{\beta})$ may be calculated by replacing $\boldsymbol{z}_{\mathcal{A}}$ and $\boldsymbol{z}$ with $\tilde{\boldsymbol{z}}_{\mathcal{A}}$ and $\tilde{\boldsymbol{z}}$, respectively, in events (7) and (8).
Remark 3 As $|\boldsymbol{\beta}_{j}^{*}|\to\infty$ for all $j\in\mathcal{A}$, the probability of sign recovery and support recovery are equivalent. To see this, note that $\textbf{Z}_{\mathcal{A}}\textbf{V}^{1/2}_{\mathcal{A}}\boldsymbol{\beta}_{\mathcal{A}}=\textbf{V}^{1/2}_{\mathcal{A}}|\boldsymbol{\beta}_{\mathcal{A}}|$ so $P(S_{\lambda}\,|\,\textbf{F}_{\mathcal{A}},\,\boldsymbol{\beta}_{\mathcal{A}})\to 1$. For any other $\tilde{\boldsymbol{z}}\in\mathcal{Z}_{\mathcal{A}}$, $\tilde{\textbf{Z}}_{\mathcal{A}}\textbf{V}^{1/2}_{\mathcal{A}}\boldsymbol{\beta}_{\mathcal{A}}$ would have negative elements, causing $P(S_{\lambda}\,|\,\textbf{F}_{\mathcal{A}},\,\boldsymbol{\beta}_{\mathcal{A}})\to 0$. Finally, as $I_{\lambda}$ is independent of $|\boldsymbol{\beta}_{\mathcal{A}}|$, $P(\hat{\mathcal{A}}=\mathcal{A}\,|\,\textbf{F},\,\boldsymbol{\beta})\to P(\hat{\boldsymbol{z}}=\boldsymbol{z}\,|\,\textbf{F},\,\boldsymbol{\beta})$.
@@ 2.1 Local criteria assuming known $\boldsymbol{\beta}$
$\phi_{\lambda}(\textbf{X}\,|\,\boldsymbol{\beta})=P(\hat{\boldsymbol{z}}=\boldsymbol{z}\,|\,\textbf{F},\,\boldsymbol{\beta})=P(S_{\lambda}\cap I_{\lambda}\,|\,\textbf{F},\,\boldsymbol{\beta})\ .$ (9)
@@
For a given X and its F, consider a $\boldsymbol{\beta}$ and its reflection $-\boldsymbol{\beta}$. Then
$\phi_{\lambda}(\textbf{X}\,|\,\boldsymbol{\beta})=P(S_{\lambda}\,|\,\textbf{F}_{\mathcal{A}},\,\boldsymbol{\beta}_{\mathcal{A}})\times P(I_{\lambda}\,|\,\textbf{F},\boldsymbol{z})=\phi_{\lambda}(\textbf{X}\,|\,-\boldsymbol{\beta})\ ,$
and an $\textbf{X}^{*}$ optimal for $\phi_{\lambda}$ under $\boldsymbol{\beta}$ is also optimal for $\phi_{\lambda}$ under $-\boldsymbol{\beta}$.
Lemma 1 establishes events $S_{\lambda}$ and $I_{\lambda}$ are independent and provides a design equivalence result. The following equivalence theorem is more general and further simplifies the implementation of our framework:
@@
Let $\textbf{X}^{*}$ be a locally optimal design for $\phi_{\lambda}$ under a $\boldsymbol{\beta}$ where $\boldsymbol{z}_{\mathcal{A}}=\boldsymbol{1}$. Consider an alternative $\tilde{\boldsymbol{\beta}}=\tilde{\textbf{Z}}\boldsymbol{\beta}$ where $\tilde{\textbf{Z}}$ is any diagonal matrix comprised of ±$\pm 1$. Then $\phi_{\lambda}(\textbf{X}\,|\,\tilde{\boldsymbol{\beta}})=\phi_{\lambda}(\textbf{X}\tilde{\textbf{Z}}\,|\,\boldsymbol{\beta})$ and $\tilde{\textbf{X}^{*}}=\textbf{X}^{*}\tilde{\textbf{Z}}$ is locally optimal for $\phi_{\lambda}$ under $\tilde{\boldsymbol{\beta}}.$
@@
The following criterion treats $\boldsymbol{z}_{\mathcal{A}}$ as an unknown quantity:
@@
It is the expected probability of sign recovery assuming all $\tilde{\boldsymbol{z}}\in\mathcal{Z}_{\mathcal{A}}$ are equally likely. Calculating all $2^{k}$ probabilities can be computationally intensive, but Lemma 1 allows us to halve the number of computations. We state this as a corollary:
@@
For a known $\mathcal{A}$, there exists a local optimal design $\textbf{X}^{*}$ where $(\textbf{I}-\textbf{P}_{1})\textbf{X}^{*}_{\mathcal{I}}=\boldsymbol{0}$, making $\phi_{\lambda}(\textbf{X}^{*}\,|\,\boldsymbol{\beta})=P(S_{\lambda}\,|\,\textbf{F}_{\mathcal{A}},\boldsymbol{\beta}_{\mathcal{A}})$ and $\phi_{\lambda}^{\pm}(\textbf{X}^{*}\,|\,\boldsymbol{\beta})=2^{-(k-1)}\sum_{\tilde{z}\in\mathcal{Z}^{\pm}_{\mathcal{A}}}P(S_{\lambda}\,|\,\textbf{F}_{\mathcal{A}}\tilde{\textbf{Z}},\boldsymbol{\beta}_{\mathcal{A}})$.
Local optimal designs exploit knowledge about $\mathcal{A}$ by completely confounding the columns of the inactive factors with the intercept, making $P(I_{\lambda}\,|\,\textbf{F},\,\boldsymbol{\beta})=1$. Finding the local optimal design, which was originally supersaturated, then reduces to finding a full column rank matrix $\textbf{X}_{\mathcal{A}}$ that maximizes probability of the $S_{\lambda}$ event(s).
To demonstrate this new theory, we will now show that if $\boldsymbol{z}_{\mathcal{A}}=\boldsymbol{1}$ is assumed known, an orthogonal design may not maximize $P(S_{\lambda}\,|\,\textbf{F}_{\mathcal{A}},\boldsymbol{\beta}_{\mathcal{A}})$. For any $\textbf{X}_{\mathcal{A}}$ the bounds of integration for $S_{\lambda}$ and the mean of $\boldsymbol{u}$ are proportional to $\textbf{V}_{\mathcal{A}}^{1/2}\boldsymbol{\beta}_{\mathcal{A}}$ and $\lambda\textbf{C}_{\mathcal{A}}^{-1}\boldsymbol{1}$, respectively. Hence, a design that maximizes $\textbf{V}_{\mathcal{A}}^{1/2}\boldsymbol{\beta}_{\mathcal{A}}-\lambda\textbf{C}_{\mathcal{A}}^{-1}\boldsymbol{1}$ may maximize $P(S_{\lambda}\,|\,\textbf{F}_{\mathcal{A}},\boldsymbol{\beta}_{\mathcal{A}})$, although the covariance matrix of $\boldsymbol{u}$ also plays a role. For any orthogonal design, this difference is $\boldsymbol{\beta}_{\mathcal{A}}-\lambda\boldsymbol{1}$. We will try to find a $\textbf{X}_{\mathcal{A}}$ with larger sign recovery probability by identifying a design whose $\textbf{V}_{\mathcal{A}}^{1/2}$ and $\textbf{C}_{\mathcal{A}}$ satisfy
@@
with equality if and only if $k_{2}=n/2$. A similar expression holds for $\xi_{2}$ with $\tilde{k}_{1}$ in the numerator of (15), and $\xi_{2}=0$ if and only if $k_{1}=n/2$. Since the diagonal elements of $\textbf{C}_{\mathcal{A}}$ are 1, $\rho_{ii}\geq 1$ and it follows $\boldsymbol{1}-\boldsymbol{\xi}>\boldsymbol{0}$. Hence designs constructed in this way satisfy (11) for some combinations of $\boldsymbol{\beta}_{\mathcal{A}}$ and $\lambda$, but whether this implies higher sign recovery probability depends on the covariance matrix of $\boldsymbol{u}$.
For example, suppose $n=16$ and we have $p=20$ factors but we know the first $k=8$ factors are nonzero. We considered three SSDs that set all the inactive factors columns to $+1$. The first SSD set the $8$ active factor columns to an orthogonal $\textbf{X}_{\mathcal{A}}$ and the other two SSDs set active effects columns according to (14) with $k_{1}=4$ and $8$, respectively. For all designs, $\textbf{C}_{\mathcal{A}}^{-1}\boldsymbol{1}=\xi\boldsymbol{1}$ where $\xi=1$, $0.1575$, and $0.1607$ for the orthogonal, $k_{1}=4$, and $k_{1}=8$ designs, respectively. For the $k_{1}=4$ and $k_{1}=8$ designs, (11) is equivalent to $\lambda^{-1}\boldsymbol{\beta}_{\mathcal{A}}\leq 53.92\times\boldsymbol{1}$ and $\lambda^{-1}\boldsymbol{\beta}_{\mathcal{A}}\leq 53.72\times\boldsymbol{1}$, respectively. The three designs’ $\phi_{\lambda}$ and $\phi_{\lambda}^{\pm}$ values were compared across a range of $\lambda$ and three scenarios for $\boldsymbol{\beta}_{\mathcal{A}}$: $(0.3,0.4,\dots,1)^{T}$, $\boldsymbol{1}$, and $3\times\boldsymbol{1}$. The results for $\phi_{\lambda}$ are shown in the left panels of Figure 1.
@@
For all scenarios, there is a range of large $\lambda$ values where all designs have $\phi_{\lambda}=0$, a middle range where the two proposed designs outperform the orthogonal design, and a range of small $\lambda$ values where the orthogonal design is superior. The orthogonal design improves over the other two designs well before condition (11) is violated, due to the role of the covariance matrices. The improvement is negligible when $\boldsymbol{\beta}_{\mathcal{A}}=3\times\boldsymbol{1}$, implying the covariance matrix is less important as the elements of $\boldsymbol{\beta}_{\mathcal{A}}$ increases. The results for $\phi_{\lambda}^{\pm}$ are shown in the right panels of Figure 1, and clearly favor the orthogonal design.
@@ 2.2 Relaxed local criteria
To make the criteria more practical, we introduce relaxations regarding the choice of $\lambda$ and the assumptions of the underlying model. To reduce the dependency on $\lambda$, we propose two summary measures of $\phi_{\lambda}$ and $\phi_{\lambda}^{\pm}$. The first summary takes the maximum sign recovery probability: $\phi_{\text{max}}(\textbf{X}\,|\,\boldsymbol{\beta})=\max_{\lambda>0}\phi_{\lambda}(\textbf{X}\,|\,\boldsymbol{\beta})$. However, in practice one will perform some tuning parameter selection strategy to choose a $\lambda$, and so the benefits of a design that optimizes $\phi_{\text{max}}$ will only be realized if the strategy reliably picks the corresponding $\lambda$ that maximizes $\phi_{\lambda}$. Hence, we would like a design that maintains large probabilities for a wide range of $\lambda$. It is common to calculate the lasso solution path with respect to $\log(\lambda)$. Such a transformation stretches the region of $\lambda\in(0,1)$, whose corresponding lasso estimates would receive the least amount of penalization. Therefore, we propose the criterion
@@
The definitions for $\phi_{\max}^{\pm}(\textbf{X}\,|\,\boldsymbol{\beta})$ and $\phi_{\Lambda}^{\pm}(\textbf{X}\,|\,\boldsymbol{\beta})$ relax assumptions about $\boldsymbol{z}_{\mathcal{A}}$ and measure the maximum sign recovery and the integrated average sign recovery.
Turning to relaxing assumptions about the $\boldsymbol{\beta}$, we first assume $|\boldsymbol{\beta}_{\mathcal{A}}|\geq\beta\boldsymbol{1}$ where $\beta$ is the minimum practically significant effect size. Such a $\beta$ value is commonly chosen while performing power calculations so we will calculate our support recovery probability under the most challenging scenario of $|\boldsymbol{\beta}_{\mathcal{A}}|=\beta\boldsymbol{1}$. We will hold a discussion about incorporating uncertainty about $|\boldsymbol{\beta}_{\mathcal{A}}|$ to Section 6. To address uncertainty about $\mathcal{A}$, we fix $k\leq n-1$ and assume all supports of this size are equally likely. Denote this set of supports by $\mathcal{A}_{k}$.
With the above relaxed assumptions, we first introduce a criterion that incorporates uncertainty about $\mathcal{A}$, but treats $\boldsymbol{z}_{\mathcal{A}}$ as known. This resembles $\boldsymbol{\beta}$’s where the $Var(s+)$-optimal designs performed best in previous simulation studies. Suppose $z_{j}$ depends only on whether $j\in\mathcal{A}$ or $\mathcal{I}$, and let $\boldsymbol{z}^{*}$ be the $p\times 1$ vector comprised of the elements $z_{j}$ assuming $j\in\mathcal{A}$. Following Theorem 1, we assume without loss of generality that $\boldsymbol{z}^{*}=\boldsymbol{1}$ to identify the optimal design and then multiply the columns of the design by their actual corresponding $z_{j}^{*}$. For a given $\mathcal{A}$, let A be the $p\times p$ diagonal matrix of all zeroes except for the diagonal elements corresponding to $\mathcal{A}$, which are set to 1. Then the true sign vector is $\boldsymbol{z}=\textbf{A}\boldsymbol{1}$ and the sign-dependent criterion for a given $\beta$ is

### v3

@@ 2 Exact Local Optimality Criteria
Remark 2 The probability of support recovery, $P(\hat{\mathcal{A}}=\mathcal{A}\,|\,\textbf{X},\,\boldsymbol{\beta},\,\lambda)$, requires consideration of all possible $2^{k}$ sign vectors that have the same 00 elements as $\boldsymbol{z}$ but alternative ±$\pm 1$ elements indexed by $\mathcal{A}$. Defining $\mathcal{Z}_{\mathcal{A}}$ as the set of all such sign vectors, we have $P(\hat{\mathcal{A}}=\mathcal{A}\,|\,\textbf{X},\,\boldsymbol{\beta},\,\lambda)=2^{-k}\sum_{\tilde{\boldsymbol{z}}\in\mathcal{Z}_{\mathcal{A}}}P(\hat{\boldsymbol{z}}=\tilde{\boldsymbol{z}}\,|\,\textbf{X},\,\boldsymbol{\beta},\,\lambda)$. Each $P(\hat{\boldsymbol{z}}=\tilde{\boldsymbol{z}}\,|\,\textbf{X},\,\boldsymbol{\beta},\,\lambda)$ may be calculated by replacing $\boldsymbol{z}_{\mathcal{A}}$ and $\boldsymbol{z}$ with $\tilde{\boldsymbol{z}}_{\mathcal{A}}$ and $\tilde{\boldsymbol{z}}$, respectively, in events (7) and (8).
Remark 3 As $|\beta_{j}^{*}|\to\infty$ for all $j\in\mathcal{A}$, the probability of sign recovery and support recovery are equivalent. To see this, note that $\textbf{Z}_{\mathcal{A}}\textbf{V}^{1/2}_{\mathcal{A}}\boldsymbol{\beta}_{\mathcal{A}}=\textbf{V}^{1/2}_{\mathcal{A}}|\boldsymbol{\beta}_{\mathcal{A}}|$ so $P(S_{\lambda}\,|\,\textbf{C}_{\mathcal{A}},\,\textbf{V}_{\mathcal{A}},\,\boldsymbol{\beta}_{\mathcal{A}})\to 1$ since the area of integration approaches $\mathbb{R}^{k}$, the support of $\boldsymbol{u}$. For any other $\tilde{\boldsymbol{z}}\in\mathcal{Z}_{\mathcal{A}}$, $\tilde{\textbf{Z}}_{\mathcal{A}}\textbf{V}^{1/2}_{\mathcal{A}}\boldsymbol{\beta}_{\mathcal{A}}$ would have negative elements, causing $P(S_{\lambda}\,|\,\textbf{C}_{\mathcal{A}},\,\textbf{V}_{\mathcal{A}},\,\boldsymbol{\beta}_{\mathcal{A}})\to 0$ because the region of integration for the $u_{j}$ corresponding to these negative elements approaches $(-\infty,\lim_{t\to-\infty}t)$, which has zero probability. Hence $P(\hat{\boldsymbol{z}}=\tilde{\boldsymbol{z}}\,|\,\textbf{X},\,\boldsymbol{\beta},\,\lambda)\to 0$ for any $\tilde{\boldsymbol{z}}\neq\boldsymbol{z}$. Finally, $I_{\lambda}$ depends on $\boldsymbol{\beta}_{\mathcal{A}}$ only through $\boldsymbol{z}_{\mathcal{A}}$ so it is unaffected by increasing $|\beta_{j}^{*}|$, hence $P(\hat{\mathcal{A}}=\mathcal{A}\,|\,\textbf{X},\,\boldsymbol{\beta})\to P(\hat{\boldsymbol{z}}=\boldsymbol{z}\,|\,\textbf{X},\,\boldsymbol{\beta})$.
@@ 2.1 Local criteria assuming known $\boldsymbol{\beta}$
$\phi_{\lambda}(\textbf{X}\,|\,\boldsymbol{\beta})=P(\hat{\boldsymbol{z}}=\boldsymbol{z}\,|\,\textbf{X},\,\boldsymbol{\beta})=P(S_{\lambda}\cap I_{\lambda}\,|\,\textbf{C},\,\textbf{V}_{\mathcal{A}},\,\boldsymbol{\beta})\ .$ (9)
@@
For a given X, events $S_{\lambda}$ and $I_{\lambda}$ are independent, and $\phi_{\lambda}(\textbf{X}\,|\,\boldsymbol{\beta})=\phi_{\lambda}(\textbf{X}\,|\,-\boldsymbol{\beta})$ for any $\boldsymbol{\beta}$ and its reflection, $-\boldsymbol{\beta}$. Hence an optimal $\textbf{X}^{*}$ for $\phi_{\lambda}$ under $\boldsymbol{\beta}$ is also optimal for $\phi_{\lambda}$ under $-\boldsymbol{\beta}$.
Lemma 1 includes a design equivalence result between $\boldsymbol{\beta}$ and its reflection, $-\boldsymbol{\beta}$. The following equivalence theorem is more general and further simplifies the implementation of the framework:
@@
For a $\boldsymbol{\beta}$ where $\boldsymbol{z}_{\mathcal{A}}=\boldsymbol{1}$, consider $\tilde{\boldsymbol{\beta}}=\tilde{\textbf{Z}}\boldsymbol{\beta}$ where $\tilde{\textbf{Z}}$ is any diagonal matrix comprised of ±$\pm 1$. Then for any X, $\phi_{\lambda}(\textbf{X}\,|\,\boldsymbol{\beta})=\phi_{\lambda}(\textbf{X}\tilde{\textbf{Z}}\,|\,\tilde{\boldsymbol{\beta}})$. Consequently, if $\textbf{X}^{*}$ is optimal for $\phi_{\lambda}$ under a $\boldsymbol{\beta}$, then $\tilde{\textbf{X}^{*}}=\textbf{X}^{*}\tilde{\textbf{Z}}$ is optimal for $\phi_{\lambda}$ under $\tilde{\boldsymbol{\beta}}.$
@@
The following criterion assumes the magnitudes of the $\beta_{j}$ are known and treats $\boldsymbol{z}_{\mathcal{A}}$ as an unknown quantity and calculates the expected probability of sign recovery assuming all $\tilde{\boldsymbol{z}}\in\mathcal{Z}_{\mathcal{A}}$ are equally likely. Without loss of generality, the magnitudes are expressed by $\boldsymbol{\beta}>\boldsymbol{0}$ and the criterion is defined to be
@@
Calculating all $2^{k}$ probabilities can be computationally intensive, but Lemma 1 allows us to halve the number of computations. We state this as a corollary:
@@
For a known $\mathcal{A}$, there exists an optimal design $\textbf{X}^{*}$ where $\textbf{X}^{*}_{\mathcal{I}}=\textbf{J}$. Furthermore, the criteria $\phi_{\lambda}(\textbf{X}^{*}\,|\,\boldsymbol{\beta})$ and $\phi_{\lambda}^{\pm}(\textbf{X}^{*}\,|\,\boldsymbol{\beta})$ for such designs depend only on $S_{\lambda}$.
Optimal designs for criteria with perfect knowledge about $\mathcal{A}$ completely confound the columns of the inactive factors with the intercept, making $P(I_{\lambda}\,|\,\textbf{C},\,\boldsymbol{z})=1$. Finding the optimal design, which may originally have been supersaturated, then reduces to finding a full column rank matrix $\textbf{X}_{\mathcal{A}}$ that maximizes probability of the $S_{\lambda}$ event(s).
To demonstrate this new theory, we will now show that if $\boldsymbol{z}_{\mathcal{A}}=\boldsymbol{1}$ is assumed known, an orthogonal design may not maximize $P(S_{\lambda}\,|\,\textbf{C}_{\mathcal{A}},\,\textbf{V}_{\mathcal{A}},\,\boldsymbol{\beta}_{\mathcal{A}})$. For any $\textbf{X}_{\mathcal{A}}$ the bounds of integration for $S_{\lambda}$ and the mean of $\boldsymbol{u}$ are proportional to $\textbf{V}_{\mathcal{A}}^{1/2}\boldsymbol{\beta}_{\mathcal{A}}$ and $\lambda\textbf{C}_{\mathcal{A}}^{-1}\boldsymbol{1}$, respectively. Hence, a design that maximizes $\textbf{V}_{\mathcal{A}}^{1/2}\boldsymbol{\beta}_{\mathcal{A}}-\lambda\textbf{C}_{\mathcal{A}}^{-1}\boldsymbol{1}$ may maximize $P(S_{\lambda}\,|\,\textbf{C}_{\mathcal{A}},\,\textbf{V}_{\mathcal{A}},\,\boldsymbol{\beta}_{\mathcal{A}})$, although the covariance matrix of $\boldsymbol{u}$ also plays a role. For any orthogonal design, this difference is $\boldsymbol{\beta}_{\mathcal{A}}-\lambda\boldsymbol{1}$. We will try to find a $\textbf{X}_{\mathcal{A}}$ with larger $P(S_{\lambda}\,|\,\textbf{C}_{\mathcal{A}},\,\textbf{V}_{\mathcal{A}},\,\boldsymbol{\beta}_{\mathcal{A}})$ by identifying a design whose $\textbf{V}_{\mathcal{A}}^{1/2}$ and $\textbf{C}_{\mathcal{A}}$ satisfy
@@
with equality if and only if $k_{2}=n/2$. A similar expression holds for $\xi_{2}$ with $\tilde{k}_{1}$ in the numerator of (15), and $\xi_{2}=0$ if and only if $k_{1}=n/2$. Since the diagonal elements of $\textbf{C}_{\mathcal{A}}$ are 1, $\rho_{ii}\geq 1$ and it follows $\boldsymbol{1}-\boldsymbol{\xi}>\boldsymbol{0}$. Hence designs constructed in this way satisfy (11) for some combinations of $\boldsymbol{\beta}_{\mathcal{A}}$ and $\lambda$, but whether this implies higher $P(S_{\lambda}\,|\,\textbf{C}_{\mathcal{A}},\,\textbf{V}_{\mathcal{A}},\,\boldsymbol{\beta}_{\mathcal{A}})$ depends on the covariance matrix of $\boldsymbol{u}$.
For example, suppose $n=16$ and we have $p=20$ factors but we know the first $k=8$ factors have nonzero effects and the rest are inactive. We considered three SSDs that set all inactive factors’ columns to $+1$, eliminating any ability to estimate them. The first SSD set the $8$ active factor columns to an orthogonal $\textbf{X}_{\mathcal{A}}$ and the other two SSDs set active effects columns according to (14) with $k_{1}=4$ and $8$, respectively. For all designs, $\textbf{C}_{\mathcal{A}}^{-1}\boldsymbol{1}=\xi\boldsymbol{1}$ where $\xi=1$, $0.1575$, and $0.1607$ for the orthogonal, $k_{1}=4$, and $k_{1}=8$ designs, respectively. For the $k_{1}=4$ and $k_{1}=8$ designs, (11) is equivalent to $\lambda^{-1}\boldsymbol{\beta}_{\mathcal{A}}\leq 53.92\times\boldsymbol{1}$ and $\lambda^{-1}\boldsymbol{\beta}_{\mathcal{A}}\leq 53.72\times\boldsymbol{1}$, respectively. The three designs’ $\phi_{\lambda}$ and $\phi_{\lambda}^{\pm}$ values were compared across a range of $\lambda$ and three scenarios for $\boldsymbol{\beta}_{\mathcal{A}}$: $(0.3,0.4,\dots,1)^{T}$, $\boldsymbol{1}$, and $3\times\boldsymbol{1}$. The results for $\phi_{\lambda}$ are shown in the left panels of Figure 1.
@@
For all scenarios, there is a range of large $\lambda$ values where all designs have $\phi_{\lambda}=0$, a middle range where the two proposed designs outperform the orthogonal design, and a range of small $\lambda$ values where the orthogonal design is superior. The orthogonal design improves over the other two designs before condition (11) is violated, due to the role of the covariance matrices. The improvement is negligible when $\boldsymbol{\beta}_{\mathcal{A}}=3\times\boldsymbol{1}$, implying the covariance matrix is less important as the elements of $\boldsymbol{\beta}_{\mathcal{A}}$ increases. The results for $\phi_{\lambda}^{\pm}$ are shown in the right panels of Figure 1, and clearly favor the orthogonal design.
@@ 2.2 Relaxed local criteria
To make the criteria in Section 2.1 more practical, we introduce relaxations regarding the choice of $\lambda$ and the assumptions of the underlying model. To reduce the dependency on $\lambda$, we propose two summary measures of $\phi_{\lambda}$ and $\phi_{\lambda}^{\pm}$. The first summary takes the maximum sign recovery probability: $\phi_{\text{max}}(\textbf{X}\,|\,\boldsymbol{\beta})=\max_{\lambda>0}\phi_{\lambda}(\textbf{X}\,|\,\boldsymbol{\beta})$. However, in practice one will perform some tuning parameter selection strategy to choose a $\lambda$, and so the benefits of a design that optimizes $\phi_{\text{max}}$ will only be realized if the strategy reliably picks the corresponding $\lambda$ that maximizes $\phi_{\lambda}$. Hence, we would like a design that maintains large probabilities for a wide range of $\lambda$. It is common to calculate the lasso solution path with respect to $\log(\lambda)$. Such a transformation stretches the region of $\lambda\in(0,1)$, whose corresponding lasso estimates would receive the least amount of penalization. Therefore, we propose the criterion
@@
The definitions for $\phi_{\max}^{\pm}(\textbf{X}\,|\,\boldsymbol{\beta})$ and $\phi_{\Lambda}^{\pm}(\textbf{X}\,|\,\boldsymbol{\beta})$ relax assumptions about $\boldsymbol{z}_{\mathcal{A}}$ and measure the maximum expected sign recovery probability and the integrated expected sign recovery probability, assuming all $\boldsymbol{z}_{\mathcal{A}}$ are equally likely.
Turning to relaxing assumptions about the $\boldsymbol{\beta}$, we first assume $|\boldsymbol{\beta}_{\mathcal{A}}|\geq\beta\boldsymbol{1}$ where $\beta$ is the minimum practically significant effect size. Such a $\beta$ value is commonly chosen while performing power calculations so we will calculate the support recovery probability under the most challenging scenario of $|\boldsymbol{\beta}_{\mathcal{A}}|=\beta\boldsymbol{1}$. We will defer the discussion about incorporating uncertainty about $|\boldsymbol{\beta}_{\mathcal{A}}|$ to Section 6. Finally, to address uncertainty about $\mathcal{A}$, we fix $k\leq n-1$ and assume all supports of this size are equally likely. Denote this set of supports by $\mathcal{A}_{k}$.
With the above relaxed assumptions, we first introduce a criterion that incorporates uncertainty about $\mathcal{A}$, but treats $\boldsymbol{z}_{\mathcal{A}}$ as known. This resembles scenarios where $\text{Var}(s+)$-optimal designs have been shown to outperform $\text{E}(s^{2})-$ and $\text{UE}(s^{2})-$optimal SSDs in simulation studies. Assume $z_{j}$ depends only on whether $j\in\mathcal{A}$ or $\mathcal{I}$ as follows. Let $\boldsymbol{z}^{*}$ be the $p\times 1$ vector comprised of ±$\pm 1$ for each $j$, assuming $j\in\mathcal{A}$. Then $z_{j}=z_{j}^{*}$ when $j\in\mathcal{A}$ and $z_{j}=0$ when $j\in\mathcal{I}$. Following Theorem 1, we assume without loss of generality that $\boldsymbol{z}^{*}=\boldsymbol{1}$ to identify the optimal design and then multiply the columns of the design by their actual corresponding $z_{j}^{*}$. For a given $\mathcal{A}$, let A be the $p\times p$ diagonal matrix of all zeroes except for the diagonal elements corresponding to $\mathcal{A}$, which are set to 1. Then the true sign vector is $\boldsymbol{z}=\textbf{A}\boldsymbol{1}$ and the sign-dependent criterion for a given $\beta$ is